Dan's biceps muscle has a moment arm about the elbow joint of 0.011 m when the joint is held at a joint angle of 90 degrees and can produce a force of 526 N. His forearm is 0.24 m from the elbow joint center to the center of his hand where he grips a handle. His forearm has a mass of 1.2 kg and the center of mass is located directly above the elbow joint. How much force can he deliver to the handle with force from his biceps?
Answer #1

moment about center of mass = 0.011*526= 5.786

since system is in eqilibrium, therefore ,

5.786= (1.2*10 + x) *0.24

24.1107 = 12 + x

x =12.1107 N